# Enhancing System-Level Safety in Mixed-Autonomy Platoon via Safe Reinforcement Learning

### 20 Jan 2024 v1

@@ Abstract
Connected and automated vehicles (CAVs) have recently gained prominence in traffic research, thanks to the advancements in communication technology and autonomous driving. A variety of longitudinal control strategies for CAVs have been developed to enhance traffic efficiency, stability, and safety in mixed-autonomy scenarios. Deep reinforcement learning (DRL) is one promising strategy for mixed-autonomy platoon control since it can tackle complex scenarios in real-time. However, there are three research gaps for DRL-based mixed-autonomy platoon control. First, incorporating safety considerations into DRL typically relies on designing collision avoidance-based reward functions, which lack collision-free guarantees. Second, current DRL-based-control approaches for mixed traffic only consider the safety of CAVs, with little attention paid to the surrounding HDVs. To address the research gaps, we introduce a differentiable safety layer that converts DRL actions into safe actions with collision-free guarantees. This process relies on solving a differentiable quadratic programming problem that incorporates control barrier function-based (CBF) safety constraints for both CAV and its following HDVs to achieve system-level safety. Moreover, constructing CBF constraints needs system dynamics for the following HDVs, and thus we employ an online system identification module to estimate the car-following dynamics of the surrounding HDVs. The proposed safe reinforcement learning approach explicitly integrates system-level safety constraints into the training process and enables our method to adapt to varying safety-critical scenarios. Simulation results demonstrate that our proposed method effectively ensures CAV safety and improves HDV safety in mixed platoon environments while simultaneously enhancing traffic capacity and string stability.
@@ I Introduction
Connected and automated vehicles (CAVs) have been widely recognized to be beneficial for traffic flow, thanks to their capability of exchanging real-time information and eliminating undesirable human driving behavior [1, 2, 3, 4]. Significant research efforts have been placed on longitudinal control, i.e., the coordinated control of the speed profile of a platoon of CAVs, with enormous potential to enhance road capacity [5], traffic stability [6, 7], energy efficiency [8, 9, 10], and safety [11, 4, 12]. Early research on longitudinal control tends to assume all vehicles to be CAVs [13]. However, such an assumption is not practical in the near future since the penetration rates of CAVs can only increase gradually as the technology matures and public acceptance improves. Therefore, it is crucial to investigate mixed-autonomy platoons whereby CAVs and human-driven vehicles (HDVs) coexist.
@@
Several control schemes have been proposed in the context of mixed-autonomy platoons, whereby CAVs are controlled considering the behavior of HDVs around them. One typical scheme is connected cruise control [14], whereby the CAV at the tail of a platoon adjusts its driving decisions considering multiple HDVs ahead. However, such a scheme only considers the preceding vehicles of the studied CAV without accounting for the CAV’s impact on the following vehicles. In contrast, the recently proposed scheme of leading cruise control (LCC) extends traditional cruise control by allowing CAVs to utilize information from both preceding and following HDVs [6]. With the additional information and flexibility compared to connected cruise control, LCC offers a better potential to improve the performance of the entire platoon.
@@
Various controllers have been developed to implement LCC, including linear feedback controllers [6], optimization-based controllers using model predictive control (MPC) [15, 16] and data-enabled predictive control (DeePC) [17, 18], and learning-based controllers based on deep reinforcement learning (DRL) [19, 20]. Linear feedback controllers are easy to design but can hardly handle complex constraints and optimization objectives. MPC can explicitly incorporate complex constraints and objectives into an embedded optimization problem, which, however, relies on explicit system modeling and requires accurate system identification to estimate system parameters. Although DeePC can simultaneously perform system identification and control in a data-driven non-parametric manner, it may require a large amount of computational resources to update the decision variables in each decision step. Among all these controllers, DRL-based controllers appear promising in that they can handle complex scenarios with a marginal online computational burden.
@@
However, existing research on DRL-based mixed-autonomy platoon control suffers from three limitations. First, safety is only indirectly considered by incorporating safety-related penalties into the reward function, which can hardly provide safety guarantees due to the black-box nature of RL. Second, existing works tend to assume skilled and rational human drivers who will not collide with their preceding vehicles, and consequently only consider ego-vehicle safety [21, 22], i.e., the safety spacing between the controlled CAV and its preceding vehicle. However, such an assumption is not realistic since human errors are common. Undesirable behavior of the following HDVs may place other vehicles in the platoon in dangerous situations and thereby undermine system-level safety for both HDVs and CAVs. Moreover, we recognize that CAVs, once detecting such undesirable behavior of HDVs, have the potential to adjust their acceleration to improve the safety performance of the entire platoon. Therefore, it is beneficial to consider system-level safety when controlling CAVs. Third, existing literature often assumes the human driver model to be known to CAVs, which, nevertheless, does not hold in real traffic systems, where driver behavior can be time-varying and diverse.
@@
To address these limitations, we propose a safe DRL-based controller that can provide a system-level safety guarantee for mixed-autonomy platoon control. Our work extends [23] that incorporated safety guarantees into a supervised learning-based controller via a quadratic programming (QP)-based safety filter, whereby the safety constraints are characterized by control barrier functions (CBF) [24]. It is worth noting that the canonical form of [23] is not capable of addressing the aforementioned limitations, and hence, we extend [23] from the following three perspectives. First, [23] leveraged supervised learning to imitate an optimal controller whose decisions are pre-calculated and used as labels, which does not fit our case where labels are unavailable. We address this limitation by devising a differentiable safety layer for the DRL-based controller in an unsupervised manner to provide safety guarantees in both online training and offline testing processes. Second, applying CBF to protect system-level safety requires the use of high-order CBFs [25], which involves finding high-order derivatives and can be computationally expensive and imprecise in real systems. To address this issue, we introduce reduced order CBF candidates to make the relative degree for CBF candidates of the following vehicles remain $1$ to avoid calculating high-order derivatives. Third, note that constructing CBF requires the explicit modeling of HDVs’ system dynamics, which is nevertheless unknown in the real traffic scenario. To address the issue, we employ a learning-based human driver behavior identification approach to characterize the unknown car-following behavior.
@@
A preliminary version of this paper was presented at the 2023 Intelligent Transportation Systems Conference. In this full version, we have made the following extensions: (i) generalizing the proposed methods from considering solely CAV safety to encompass system-level safety for both CAVs and HDVs to handle potentially undesirable HDV behavior, (ii) introducing an online learning-based system identification approach to estimate the unknown car-following dynamics of surrounding HDVs, and (iii) presenting new simulation results, e.g., safety regions for both CAVs and HDVs, to better illustrate the efficacy of our proposed method in enhancing safety performance.
@@ II-A Reinforcement Learning (RL)
The function $\text{clip}(\cdot)$ prevents aggressive updating by utilizing a trust region defined by a hyperparameter $\epsilon$. $A^{\pi_{\theta_{\text{RL},\text{old }}}}(\bm{x}_{t},\bm{u}_{t})$ is the advantage function evaluating the benefit of selected actions.
@@ III Mixed-Autonomy Traffic Environment Modeling
We consider a mixed-autonomy platoon comprising HDVs (denoted by set $\Omega_{\mathcal{H}}$) and CAVs (denoted by set $\Omega_{\mathcal{C}}$), whereby the driving behavior of HDVs is modeled using car-following models (unknown to CAVs) and CAVs are controlled by RL-based controllers. CAV $i\in\Omega_{\mathcal{C}}$ collects state information within the perception range from both preceding and following vehicles in sets $\Omega_{\mathcal{P},i}$ and $\Omega_{\mathcal{F},i}$, respectively, and determine its control action based on the collected information. Moreover, we envision that the CAV penetration rate will remain very low at the early stage of CAV deployment, and hence it would be common for one CAV to be surrounded by multiple HDVs. Therefore, we assume that only one CAV presents in the platoon, i.e., $|\Omega_{\mathcal{C}}|=1$ with ||$\cdot$|| ⋅ | denotes the cardinality of a set.
For each vehicle $i\in\Omega_{\mathcal{C}}\cup\Omega_{\mathcal{H}}$, we consider dynamics described by second-order ordinary differentiable equations with states including the speed of vehicle $i$ (i.e., $v_{i}$) and the spacing between vehicle $i$ and vehicle $i-1$ (i.e., $s_{i}$).
@@
$\left\{\begin{array}[]{l}\dot{s}_{i}(t)=v_{i-1}(t)-v_{i}(t),\\ \dot{v}_{i}(t)=u_{i}(t).\end{array}\quad i\in\Omega_{\mathcal{C}},\right.$ (5)
@@
where the acceleration rate is determined by a car-following model $\mathbb{F}(\cdot)$ as a function of the spacing $s_{i}(t)$, the velocity of the preceding vehicle $v_{i-1}(t)$, and its own velocity $v_{i}(t)$.
Without loss of generality, we choose the optimal velocity model (OVM) [28] as the car-following model $\mathbb{F}$, following the setting of LCC [6]. We make the following two remarks. First, unlike [6] that assumes the model to be a public knowledge, we make a realistic assumption that this model (i.e, the formulation and parameters) remains undisclosed to the CAV, and the CAV can only learn the behavior of HDVs from its observations. Second, the OVM can be replaced by any car-following model without influencing the applicability of the proposed method. The OVM reads as follows:
@@
IV Safety-Critical Learning-based Control for CAV Platoons
@@ IV-A RL-based Controller
We formulate the control of mixed-autonomy platoons as an MDP $\mathcal{M}=(\mathcal{X},\mathcal{U},P,r,\gamma)$, where $\mathcal{X}$ includes all possible states $x(t)$ defined by the spacing and velocity, $\mathcal{U}$ comprises all possible actions $u(t)$ defined by the acceleration rate of the CAV, the transition probability is defined as $P=(f,B)$ that involves the vehicle dynamics $f(x)$ and control matrix $B$, and the policy of the RL-based controller is parameterized by $\theta_{\text{RL}}$. The reward function $r$ is introduced next.
@@
$R_{i}=r_{\text{stability},i}+r_{\text{efficiency},i}+r_{\text{safety},i}$ (10)
where $R_{i}$ represents the total reward, and $r_{\text{stability},i}$, $r_{\text{efficiency},i}$, and $r_{\text{safety},i}$ denote the string stability reward, efficiency reward, and safety reward for the CAV at position $i$, respectively. Next, these three types of rewards will be discussed in detail.
For stability, we focus on string stability [29], a desirable property of vehicle platoons such that disturbances of the lead vehicle do not get amplified when propagating through the platoon.
Specifically, we aim to minimize velocity oscillations of both the CAV and the following vehicles resulting from the disturbances of the leading vehicle $i-1$:
@@
where $\kappa_{j}\leq 1$ is the coefficient for the following vehicle velocity oscillation.
@@
$r_{\text{safety},i}=\begin{cases}\log\left(\frac{\mathrm{TTC}_{i}}{4}\right)&0\leq\mathrm{TTC}\leq 4\\ 0&\text{ otherwise }\end{cases}$ (15)
@@ IV-B Learning-Based Human Driver Behavior Identification
As in Fig. 2, we employ a learning-based human driver behavior identification module to derive the unknown car-following parameters for each HDV. The dynamics of the estimated car-following behavior for vehicle $i$ is expressed as:
$\hat{F}_{i}(s_{i},v_{i},v_{i-1})=\eta_{\xi,i}(s_{i},v_{i},v_{i-1})+\zeta_{\psi,i}(s_{i},v_{i},v_{i-1})$ (16)
where $\eta_{\xi,i}$ is a physics-based computation graph parameterized by $\xi$ that estimates a linearized car-following model, and $\zeta_{\psi,i}$ is a fully connected network parameterized by $\psi$ that estimates nonlinear bias in the car-following model. The two parts are introduced as follows.
The physics-based computation graph directly estimates the linearized car-following parameters $\xi=\left[\hat{\alpha}_{1},\hat{\alpha}_{2},\hat{\alpha}_{3}\right]$ as:
$\eta_{\xi,i}(s_{i},v_{i},v_{i-1})=\hat{\alpha}_{1}s_{i}-\hat{\alpha}_{2}v_{i}+\hat{\alpha}_{3}v_{i-1}$ (17)
where the linearized dynamics for HDVs are derived using first-order Taylor expansion as in [6].
Furthermore, the accuracy of linearized dynamics is compromised by the inherent error incurred during linearization. To address such a limitation, we augment the model with a fully connected neural network denoted as $\zeta_{\psi,i}$. Unlike classical system identification that considers only the linear components, this addition estimates the nonlinear bias in the car-following model, thereby enhancing the precision of the estimation process.
Using the estimated system dynamics for each HDV $\hat{F}_{i}$, we can formulate CBFs for both the following HDVs and the CAV.
IV-C Differentiable Safety Layer
In this subsection, we introduce the differential safety layer that comprises two main components. The first component utilizes the CBF-QP-based framework to integrate safety guarantees into the DRL-based controller by converting the DRL actions to a safey one. To take the following vehicles’ safety into account, this approach encompasses CBF-based safety constraints not only for the CAV but also for the following HDVs. Consequently, it goes beyond merely improving the safety of the ego vehicle but enhances the system-level safety of the platoon. The second component introduces a differentiable QP layer for neural networks that enables the backpropogation of the QP solution to parameters in the CBF-QP module and the DRL policy network such that both the safety layer and the DRL-based controller can be simultaneously trained. This module addresses the issue with [23] that the safety filter does not provide guidance on the training of the DRL mode.

### 01 Mar 2024 v2

@@ Abstract
Connected and automated vehicles (CAVs) have recently gained prominence in traffic research due to advances in communication technology and autonomous driving. Various longitudinal control strategies for CAVs have been developed to enhance traffic efficiency, stability, and safety in mixed-autonomy scenarios. Deep reinforcement learning (DRL) is one promising strategy for mixed-autonomy platoon control, thanks to its capability of managing complex scenarios in real time after sufficient offline training. However, there are three research gaps for DRL-based mixed-autonomy platoon control: (i) the lack of theoretical collision-free guarantees, (ii) the widely adopted but impractical assumption of skilled and rational drivers who will not collide with preceding vehicles, and (iii) the strong assumption of a known human driver model. To address these research gaps, we propose a safe DRL-based controller that can provide a system-level safety guarantee for mixed-autonomy platoon control. First, we combine control barrier function (CBF)-based safety constraints and DRL via a quadratic programming (QP)-based differentiable neural network layer to provide theoretical safety guarantees. Second, we incorporate system-level safety constraints into our proposed method to account for the safety of both CAVs and the following HDVs to address the potential collisions due to irrational human driving behavior. Third, we devise a learning-based system identification approach to estimate the unknown human car-following behavior in the real system. Simulation results demonstrate that our proposed method effectively ensures CAV safety and improves HDV safety in mixed platoon environments while simultaneously enhancing traffic capacity and string stability.
@@ I Introduction
Connected and automated vehicles (CAVs) have been widely recognized to be beneficial for traffic flow, thanks to their capability of exchanging real-time information and eliminating undesirable human driving behavior [1, 2, 3, 4, 5]. Significant research efforts have been placed on longitudinal control, i.e., the coordinated control of the speed profile of a platoon of CAVs, with enormous potential to enhance road capacity [6], traffic stability [7, 8], energy efficiency [9, 10, 11], and safety [12, 4, 13, 14, 15, 16]. Early research on longitudinal control tends to assume all vehicles to be CAVs [17, 18]. However, such an assumption is not practical in the near future since the penetration rates of CAVs can only increase gradually as the technology matures and public acceptance improves. Therefore, it is crucial to investigate mixed-autonomy platoons, whereby CAVs and human-driven vehicles (HDVs) coexist.
@@
Several control schemes have been proposed in the context of mixed-autonomy platoons, whereby CAVs are controlled considering the behavior of HDVs around them. One typical scheme is connected cruise control [19, 20], whereby the CAV at the tail of a platoon adjusts its driving decisions considering multiple HDVs ahead. However, such a scheme only considers the preceding vehicles of the studied CAV without accounting for the CAV’s impact on the following vehicles. In contrast, the recently proposed scheme of leading cruise control (LCC) extends traditional cruise control by allowing CAVs to utilize information from both preceding and following HDVs [7]. With the additional information and flexibility compared to connected cruise control, LCC offers a better potential to improve the performance of the entire platoon.
@@
Various controllers have been developed to implement LCC, including linear feedback controllers [7], optimization-based controllers using model predictive control (MPC) [21, 22] and data-enabled predictive control (DeePC) [23, 24, 25], and learning-based controllers based on deep reinforcement learning (DRL) [26, 27]. Linear feedback controllers are easy to design but can hardly handle complex constraints and optimization objectives. MPC can explicitly incorporate complex constraints and objectives into an embedded optimization problem, which, however, relies on explicit system modeling and requires accurate system identification to estimate system parameters. Although DeePC can simultaneously perform system identification and control in a data-driven nonparametric manner, it may require a large amount of computational resources to update the decision variables in each decision step. Among all these controllers, DRL-based controllers appear promising in that they can handle complex scenarios with a marginal online computational burden after sufficient offline training.
@@
However, existing research on DRL-based mixed-autonomy platoon control suffers from three limitations. First, safety is only indirectly considered by incorporating safety-related penalties into the reward function, which can hardly provide safety guarantees due to the black-box nature of RL. Second, existing works tend to assume skilled and rational human drivers who will not collide with their preceding vehicles, and consequently only consider ego-vehicle safety [28, 29], i.e., the safety spacing between the controlled CAV and its preceding vehicle. However, such an assumption is not realistic since human errors are common. The undesirable behavior of the following HDVs may place other vehicles in the platoon in dangerous situations and thereby undermine system-level safety for both HDVs and CAVs. Moreover, we recognize that CAVs, once detecting such undesirable behavior of HDVs, have the potential to adjust their acceleration to improve the safety performance of the entire platoon. Therefore, it is beneficial to consider system-level safety when controlling CAVs. Third, existing literature often assumes the human driver model to be known to CAVs, which, nevertheless, does not hold in real traffic systems, where driver behavior can be time-varying and diverse.
@@
To address these limitations, we propose a safe DRL-based controller that can provide a system-level safety guarantee for mixed-autonomy platoon control in a single-lane freeway section. Our work extends [30] that incorporated safety guarantees into a supervised learning-based controller via a quadratic programming (QP)-based safety filter, whereby the safety constraints are characterized by control barrier functions (CBF) [31]. It should be noted that the canonical form of [30] is not capable of addressing the aforementioned limitations, and hence we extend [30] from the following three perspectives. First, [30] leveraged supervised learning to imitate an optimal controller whose decisions are pre-calculated and used as labels, which does not fit our case where labels are unavailable. We address this limitation by devising a differentiable safety module for the DRL-based controller in an unsupervised manner to provide safety guarantees in both online training and offline testing processes. Second, applying CBF to protect system-level safety requires the use of high-order CBFs [32], which involves finding high-order derivatives and can be computationally expensive and imprecise in real systems. To address this issue, we introduce reduced order CBF candidates to make the relative degree for CBF candidates of the following vehicles remain $1$ to avoid calculating high-order derivatives. Third, note that constructing CBF requires the explicit modeling of HDVs’ system dynamics, which is nevertheless unknown in the real traffic scenario. To address the issue, we employ a learning-based human driver behavior identification approach to characterize the unknown car-following behavior.
@@
A preliminary version of this paper [16] was presented at the 2023 Intelligent Transportation Systems Conference. In this full version, we have made the following extensions: (i) generalizing the proposed methods from considering solely CAV safety to encompass system-level safety for both CAVs and HDVs to handle potentially undesirable HDV behavior, (ii) introducing an online learning-based system identification approach to estimate the unknown car-following dynamics of surrounding HDVs, and (iii) presenting new simulation results, e.g., safety regions for both CAVs and HDVs, to better illustrate the efficacy of our proposed method in enhancing safety performance.
@@ II-A Reinforcement Learning (RL)
The function $\text{clip}(\cdot)$ prevents aggressive updating by utilizing a trust region defined by a hyperparameter $\epsilon$. $A^{\pi_{\theta_{\text{RL},\text{old }}}}(\bm{x}_{t},\bm{u}_{t})$ is the advantage function that evaluates the benefit of selected actions.
@@ III Mixed-Autonomy Traffic Environment Modeling
We consider a mixed-autonomy platoon in one lane of a freeway segment (either straight or curved). This platoon comprises HDVs (denoted by set $\Omega_{\mathcal{H}}$) and CAVs (denoted by set $\Omega_{\mathcal{C}}$), whereby the driving behavior of HDVs is modeled using car-following models (unknown to CAVs) and CAVs are controlled by RL-based controllers. CAV $i\in\Omega_{\mathcal{C}}$ collects state information within the communication range from both preceding and following vehicles in sets $\Omega_{\mathcal{P},i}$ and $\Omega_{\mathcal{F},i}$, respectively, and determines its control action based on the collected information. Moreover, we envision that the CAV penetration rate will remain very low at the early stage of CAV deployment, and hence it would be common for a CAV to be surrounded by multiple HDVs. Therefore, we assume that only one CAV is present in the platoon, i.e., $|\Omega_{\mathcal{C}}|=1$ with ||$\cdot$|| ⋅ | denotes the cardinality of a set.
For each vehicle $i\in\Omega_{\mathcal{C}}\cup\Omega_{\mathcal{H}}$, we consider the dynamics described by second-order ordinary differentiable equations with states including the speed of vehicle $i$ (i.e., $v_{i}$) and the spacing between vehicle $i$ and vehicle $i-1$ (i.e., $s_{i}$).
@@
$\left\{\begin{array}[]{l}\dot{s}_{i}(t)=v_{i-1}(t)-v_{i}(t),\\ \dot{v}_{i}(t)=u_{i}(t),\end{array}\quad i\in\Omega_{\mathcal{C}},\right.$ (5)
@@
where the acceleration rate is determined by a car-following model $\mathbb{F}(\cdot)$ as a function of the spacing $s_{i}(t)$, the velocity of the preceding vehicle $v_{i-1}(t)$, and its own velocity $v_{i}(t)$. Note that unlike existing works [7] that assume the model $\mathbb{F}(\cdot)$ to be public knowledge, we make a realistic assumption that both the formulation and parameters of $\mathbb{F}(\cdot)$ remain unknown to the CAV, and the CAV can only learn $\mathbb{F}(\cdot)$ from its observations.
@@
IV Safety-Critical Learning-based Control for Mixed-Autonomy Platoons
@@ IV-A RL-based Controller
We formulate the control of mixed-autonomy platoons as an MDP $\mathcal{M}=(\mathcal{X},\mathcal{U},P,r,\gamma)$, where $\mathcal{X}$ includes all possible states $x(t)$ defined as the spacing and velocity, $\mathcal{U}$ comprises all possible actions $u(t)$ defined as the acceleration rate of the CAV, the transition function $P$ (i.e., system dynamics) is defined in Eq. (7) that involves the system function $f(x(t),v_{0}(t))$ and control matrix $B$, and the policy of the RL-based controller is parameterized by $\theta_{\text{RL}}$. The reward function $r$ is introduced next.
@@
$R_{i}=w_{\text{stability}}r_{\text{stability},i}+w_{\text{efficiency}}r_{\text{efficiency},i}+w_{\text{safety}}r_{\text{safety},i}$ (8)
where $R_{i}$ represents the total reward, $r_{\text{stability},i}$, $r_{\text{efficiency},i}$, and $r_{\text{safety},i}$ denote the string stability reward, efficiency reward, and safety reward for the CAV at position $i$, respectively. The weighting coefficients $w_{\text{stability}}$, $w_{\text{efficiency}}$, and $w_{\text{safety}}$ are designed by the CAV manufacturer or operator to reflect their preferences over these control objectives. Next, these three types of rewards will be discussed in detail.
For stability, we focus on string stability [35, 15], a desirable property of vehicle platoons such that disturbances of the head vehicle do not get amplified when propagating through the platoon. Specifically, we aim to minimize velocity oscillations of both the CAV and the following vehicles resulting from the disturbances of the preceding vehicle $i-1$:
@@
where $\kappa_{j}\leq 1$ is the coefficient for the following vehicle velocity oscillation. In the current implementation, we have set each weighting coefficient to 1, following [28].
@@
$r_{\text{safety},i}=\begin{cases}\log\left(\frac{\mathrm{TTC}_{i}}{4}\right),&0\leq\mathrm{TTC}\leq 4\\ 0,&\text{ otherwise }\end{cases}$ (13)
@@ IV-B Learning-Based Human Driver Behavior Identification
Inspired by [39], we employ a learning-based human driver behavior identification module to derive the unknown car-following parameters for each HDV as in Fig. 2. The dynamics of the estimated car-following behavior for vehicle $i$, denoted by $\hat{\mathbb{F}}_{i}(\cdot)$, is expressed as:
$\hat{\mathbb{F}}_{i}(s_{i},v_{i},v_{i-1})=\alpha_{1}s_{i}-\alpha_{2}v_{i}+\alpha_{3}v_{i-1}+\zeta_{\psi,i}(s_{i},v_{i},v_{i-1}),$ (14)
where $\eta_{\xi,i}(s_{i},v_{i},v_{i-1}):=\alpha_{1}s_{i}-\alpha_{2}v_{i}+\alpha_{3}v_{i-1}$ represents a physics-based computation graph parameterized by $\xi=[\alpha_{1},\alpha_{2},\alpha_{3}]$ to approximate the car-following behavior via a linearized model, and $\zeta_{\psi,i}$ is a neural network-based bias estimator parameterized by $\psi$ that quantifies the deviation between the actual nonlinear car-following dynamics and the estimations from computation graphs. The reason for dividing the car-following dynamics into a linear physics-based computation graph and a nonlinear bias estimator is that the construction of CBF requires the estimation of the partial derivatives of $\hat{\mathbb{F}}_{i}$, i.e., $\alpha_{1}=\frac{\hat{\mathbb{F}}_{i}}{\partial s_{i}}$, $\alpha_{2}=-\frac{\hat{\mathbb{{F}}_{i}}}{\partial v_{i}}$, $\alpha_{3}=\frac{\hat{\mathbb{F}}_{i}}{\partial v_{i-1}}$. Thereby, we separate the computation graph from the bias estimator to enable the explicit training of $\xi=[\alpha_{1},\alpha_{2},\alpha_{3}]$ using real data since the form and parameters of $\hat{\mathbb{F}}_{i}(\cdot)$ are not known.
It is worth noting that our system identification method is different from both (i) classical system identification considering only linear components and (ii) neural network-based car-following models. Unlike classical system identification, we estimate the nonlinear bias in the car-following model to enhance the precision of the estimation process. Unlike neural network-based car-following models that leverage a single neural network to predict acceleration rates, we combine a computation graph to explicitly estimate $\xi=[\alpha_{1},\alpha_{2},\alpha_{3}]$ to satisfy the requirement of CBF construction.
Using the estimated system dynamics for each HDV $\hat{\mathbb{F}}_{i}$, we can formulate CBFs for both the following HDVs and the CAV.
IV-C Differentiable Safety Module
In this subsection, we introduce the differential safety module that comprises two main components. The first component utilizes the CBF-QP-based framework to integrate safety guarantees into the DRL-based controller by converting the DRL actions to a safe one. To take the following vehicles’ safety into account, this approach encompasses CBF-based safety constraints not only for the CAV but also for the following HDVs. Consequently, it goes beyond merely improving the safety of the ego vehicle but enhances the system-level safety of the platoon. The second component introduces a differentiable QP layer for neural networks that enables the backpropogation of the QP solution to parameters in the CBF-QP module and the DRL policy network such that both the safety layer and the DRL-based controller can be simultaneously trained. This module addresses the issue with [30] that the safety filter does not provide guidance on the training of the DRL mode.